OpenStudy (anonymous):

Which expression is equivalent to|a| ≤3 a ≥ –3 and a ≤ 3 a ≤ –3 and a ≥ 3 a ≥ –3 or a ≤ 3 a ≤ –3 or a ≥ 3

OpenStudy (jhannybean):

|a| = \(\pm a\) Therefore it can be written \[-a \le 3\]\[a \le 3\]When you divide by a negative, you flip the sign, so \(-a \le 3 \implies a \ge -3\)

OpenStudy (jhannybean):

So what do you think your answer choice would be?